OpenStudy (kewlgeek555):

VERY EASY - Algebra I, Function Notations and Graph: Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. a. 0, -1, 1, 0, 2 b. 0, 1, 1, 2, 3 c. 0, 1, 2, 2, 3 d. 0, 1, 1, 2, 2

OpenStudy (kewlgeek555):

I need this done by today. I am really stuck on this.

OpenStudy (anonymous):

\[ f(3) = f(3-1)+f(3-2) = f(2)+f(1) = 1+0 = 1\\ f(3)=1 \]

OpenStudy (anonymous):

\[ f(4) = f(4-1)+f(4-2) = f(3)+f(2) = 1 + 1 = 2 \]

OpenStudy (anonymous):

so the first terms are \[0,1,1,2\]

OpenStudy (anonymous):

Now try to find \(f(5)\)